PQB Meeting October 8, 2015
We will meet in TL 2236 Thursday October 8 at 1:00 pm EST.
For those calling in we'll use the Blue Jeans  audio conference system.
Phone number: 1-888-240-2560 Meeting ID: 3528502673# Meeting URL: http://bluejeans.com/3528502673
- Ed Folts: Hall A beamline progress report
- Kent Paschke: Hall A Fall Beam Tests: media:NovemberPQBStudies.pdf
Notes from this meeting:
- The earliest for beam in Hall A is November 16 and the goal is 11 GeV and 85 µA.
- Bob Michaels et al.: very good progress with Hall parity DAQ and analysis and start working on Injector parity DAQ.
- Helicity board can go up to 2.5 kHz reversal. No heating problem for the Pockels Cell HV optical switch.
- Hall A Progress Report:
- Hall C double raster installed in Hall A
- Cables for cavities all run from rack to Counting House
- Still waiting on cavity receivers from John Musson
- Signal cables from halo monitors ready and available in Counting House and HV cables connected to supply
- Waiting on halo detectors and target (fork) to arrive from Virginia Tech
- Plan to install halo target before Harp work
- Halo signals are readout in scalers and should be written to EPICS channels for MCC to use
- Reinstate the EPICS readout of halo rates
- Halo detectors may not have shielding this fall
- For lumi, there maybe a need for a thick carbon solid target